Denver, CO replaced 13,000 traffic lights, each with three 150 watt incandescent bulbs, with traffic lights using 14 watt Light Emitting Diodes (LEDs). Each light has one bulb shining at any moment.(a)The original traffic lights drew how many kW from the grid? (b)The replacement traffic lights drew how many kW from the grid? (c)The replacement LED lights drew how many fewer kW from the grid compared to the incandescent lights?
Answer #1

Answer a)

Power drawn by Old Traffic Light = (13000*150)/1000 kW = 1950 k W

Answer b)

Power drawn after replacement = (13000*14)/1000 kW = 182 kW

Answer c)

Difference = 1950 - 182 = 1768 kW

Thus, the replacement LED lights drew 1768 fewer kW from the grid compared to the incandescent lights